### v1

@@ 6.6 Prompt
Discussion and Implication: Developers often encounter challenges when crafting prompts that achieve specific outcomes, manage content length, and maintain a conversational style. Prompt engineering involves feeding the model useful phrases and instructions to elicit helpful responses, which includes clearly stating the desired topic or task, and specifying constraints to guide responses within defined boundaries. Therefore, OpenAI should provide comprehensive educational resources on prompt engineering, including tutorials, webinars, and documentation. These resources should cover the principles of effective prompt design, common pitfalls, and best practices for crafting prompts that yield accurate and useful responses. Moreover, developing interactive tools that allow developers to experiment with and optimize prompts in real-time would be highly beneficial. Finally, establishing a community platforms where developers can share successful prompts, discuss challenges, and collaborate on solving common issues is crucial.
Recent advancements in prompt design (such as In-context Learning, Chain-of-Thought) offer powerful techniques for enhancing the performance of OpenAI models. These techniques involve using contextual information, reasoning through chains of logic, and augmenting responses with external data to improve the quality and relevance of the output. However, in our analysis, the number of posts discussing these advanced techniques is relatively low, indicating that developers are not paying enough attention to these advanced technologies. Therefore, OpenAI should offer extensive documentation and resources on these advanced techniques. This documentation should include detailed explanations of the underlying principles, common pitfalls, and practical examples to help developers effectively implement these techniques. Providing step-by-step guides and use case scenarios can make these resources more accessible and practical. Moreover, providing interactive tools that allow developers to experiment with and optimize these advanced techniques in real-time would be invaluable.
7 Threats to Validity
@@ 7.1 Internal Validity
Expertise in Constructing the Challenge Taxonomy: One internal threat is related to the expertise involved in constructing the challenge taxonomy. To mitigate this, individuals involved in the taxonomy construction are required to have relevant experience in OpenAI development and a foundational understanding of large language models. This ensures that the taxonomy accurately reflects the challenges faced by developers.
Determining Question Difficulty Levels: Another internal threat is the method used to determine the difficulty level of questions posed on the OpenAI forum. Our analysis shows that only 7.0% of the sampled posts had one reply marked as a “Solution”, Hence, the difficulty level is primarily determined by the number of replies each question receives rather than relying on accepted answers.
@@
Manual Labeling of Topics: We manually label the topics, which can introduce bias. To minimize this threat, two different authors independently label the topics, and a third author, a domain expert, validates these labels. Any conflicts are resolved through discussions among the authors, thereby reducing labeling bias to an acceptable minimum.
Selection of Forum Posts: Our study selectively focuses on posts within specific categories like API, Prompting, ChatGPT, and GPT builders, excluding others such as Community, Announcements, Documentation, and Forum feedback. This selection is based on the observation that excluded categories rarely discuss challenges related to OpenAI development.
7.2 External Validity
@@
Scope of Data: An external challenge arises from the temporal limitation of our data, encompassing posts up to June 2024. As OpenAI continues to develop new applications, new challenges will likely emerge, which are not captured in our current empirical study.
Incorrect initial categorization: Some developers may choose the wrong category when posting. For example, the issue is actually related to API calls, but the developer categorizes it under Prompt. To address this issue, when categorizing posts, we carefully read all the information (including questions and replies) and analyze the correct category to mitigate such errors.
7.3 Construct Validity
Accuracy of Taxonomy Categories: There is a risk that the taxonomy categories may not fully capture all relevant challenges faced by OpenAI developers. To address this, we iteratively refine the taxonomy based on feedback and consensus among experienced annotators, ensuring a comprehensive representation of challenges.
8 Related Work
8.1 OpenAI Models for Software Engineering Tasks
Recently, OpenAI models (such as GPT-3.5, GPT-4, and GPT-4o) has been successfully applied to assist various software engineering tasks, demonstrating its versatility and effectiveness in the field. Due to the limited length of the paper, we only listed some relevant studies on popular software engineering tasks.
One significant application is in code generation, where OpenAI models have been used to refine and improve code creation processes, making it easier for developers to produce high-quality code efficiently (Liu et al, 2023b, a). In program repair, OpenAI models have shown promise by automatically identifying and fixing bugs, thereby reducing the time and effort required for manual debugging (Xia and Zhang, 2023). Moreover, ChatGPT has been utilized in bug report management, streamlining the process of handling and categorizing bug reports, which enhances the overall efficiency of software maintenance (Zhang et al, 2023). In app testing, OpenAI models aid in automating the testing process, ensuring more comprehensive test coverage and faster identification of issues (Liu et al, 2023c). Additionally, OpenAI models has been employed in vulnerability detection, where it helps in identifying potential security vulnerabilities in code, contributing to the development of more secure software applications (Fu et al, 2023; Chen et al, 2023; Lu et al, 2024). Lastly, for source code summarization, ChatGPT has been used to generate concise and accurate summaries of source code, aiding developers in understanding and documenting code more effectively (Sun et al, 2023; Zhao et al, 2023a).
8.2 Studies on Developers’ Challenges
@@
In traditional software development, developers face a broad range of challenges for different application domains (such as mobile applications (Rosen and Shihab, 2016), web applications (Scoccia et al, 2021), and concurrency development (Ahmed and Bagherzadeh, 2018)). These include understanding complex requirements, managing project scope, ensuring code quality and maintainability, handling security vulnerabilities, and integrating various software components.
In the realm of deep learning, developers face distinct challenges that differ from traditional software development. These include managing and preprocessing large datasets, tuning hyperparameters, ensuring model interpretability, and addressing ethical concerns related to bias and fairness. Deployment challenges (Chen et al, 2020) are also significant, involving the integration of deep learning models into production systems, handling performance bottlenecks, and ensuring compatibility with existing infrastructure. Empirical studies have highlighted specific issues such as bugs in model optimization (Guan et al, 2023), compatibility problems (Wang et al, 2023b), and performance-related concerns that complicate the deployment and maintenance of deep learning systems (Cao et al, 2022).
Different from previous studies, we aim to fill the gap in understanding the specific challenges faced by OpenAI developers. By systematically categorizing and analyzing the posts in the OpenAI developer forum, we provide a comprehensive overview of the challenges developers encounter. Unlike previous studies that focus on broader categories or specific domains, our research delves into the unique difficulties of OpenAI development, particularly with OpenAI’s APIs and tools. Our findings offer valuable insights for both developers and the OpenAI organization, guiding future improvements in developer support and tool development. This study not only highlights common challenges but also identifies unique challenges faced by developers, providing a foundation for targeted solutions and enhancements in OpenAI’s development ecosystem.
9 Conclusion
@@
In this study, we conduct a comprehensive analysis of the challenges faced by OpenAI developers. By examining relevant posts on the OpenAI developer forum, we observe that OpenAI-related development is gaining significant traction, with developers encountering numerous challenging issues compared to traditional software development. Our goal is to analyze the underlying challenges reflected in these questions. We investigate the popularity trends on the OpenAI developer forum and the difficulty levels of the problems raised by developers. Subsequently, we manually inspect 2,364 sampled posts related to OpenAI development and construct a detailed taxonomy consisting of 6 main categories and 27 subcategories, representing the challenges that OpenAI developers encounter. Finally, we discuss the implications of our findings for various stakeholders, including developers, researchers, and the OpenAI organization.
@@
For future research, there are several avenues to explore. First, expanding the analysis to include posts from other platforms, such as GitHub and Stack Overflow, can provide a more comprehensive view of the challenges faced by OpenAI developers. Second, an in-depth study of the solutions proposed in response to the identified challenges can yield valuable insights into effective strategies and best practices. Finally, automating the categorization process using machine learning techniques can improve the efficiency and accuracy of the analysis.
@@ Acknowledgements.
Data Availability Statements
We share our collected posts, analysis scripts, and taxonomy of challenges in a GitHub repository (https://github.com/judeomg/OADF) for open science.
Declarations
Conflict of Interest The authors declare that they have no conflict of interest.

### v3

@@ 6.6. Prompt
Discussion and Implication: In recent software engineering studies, advanced prompt design techniques like Chain-of-Thought and in-context learning (Sun et al., 2024; Lu et al., 2023; Zhao et al., 2024) have been shown to significantly improve the performance of software engineering tasks when using LLMs. However, from our analysis of the OpenAI developer forum, we find that these advanced techniques have not been sufficiently utilized by developers (as evident from the relatively few questions concerning methods like Chain-of-Thought), indicating low usage. This finding suggests a gap in developer awareness or the perceived usability of these advanced prompt design techniques. We believe this observation has broader implications for the software engineering community beyond OpenAI. For example, in addition to existing API documentation, the SE community should offer more detailed use cases with online courses, or technical tutorials, particularly for advanced prompt engineering techniques. Then developing tools that assist developers in automatically generating or optimizing prompts, or in evaluating different prompt strategies across tasks, could greatly enhance their ability to leverage advanced prompting techniques. Finally, the OpenAI forum should encourage developers to share best practices and lessons learned in using advanced prompting techniques.
7. Discussions
7.1. Generalization of Our Constructed Category
To validate the generalization of our methodology, we further analyzed LLM-related issues on GitHub to construct the taxonomy. We show the process as follows. First, we focus on the official GitHub accounts of three LLM vendors: Meta AI [^86] https://www.meta.ai, Google [^87] https://www.google.com, and OpenAI. The representative LLMs of these three vendors are Llama, Gemini, and GPT. We then select projects hosted on their homepages based on the following criteria: (1) relevance to LLMs, and (2) more than 50 issues. After determining the projects, we download the issues from these specific repositories via GitMiner [^88] https://github.com/UnkL4b/GitMiner. Specifically, the data gathered for each issue includes its URL, title, submission date, status, labels, and issue description. In summary, we obtain 1,948 issues related to Llama, 560 issues related to Gemini, and 2,317 issues related to GPT. Second, we randomly sample 356 posts from these 4,825 issues with a 95% confidence level and $\pm$5confidence interval (Aghajani et al., 2019). Then we manually analyze these sampled issues by using the open coding procedure shown in Section 3. Finally, we construct a corresponding taxonomy shown in Fig. 19.
Figure 19. Challenge Taxonomy for LLM developers Built from GitHub Issues.
In general, the taxonomy constructed based on the OpenAI forum covers the categories identified in the taxonomy built from GitHub. This suggests that the taxonomy developed from the OpenAI forum is representative to some extent. Through the analysis of these two taxonomies, we identify some differences between them. Specifically, we find that in the taxonomy built from GitHub, the [A.3] Feature Suggestions category accounts for 11.7%, significantly higher than the 2.9% in the OpenAI developer forum. We conjecture that since GitHub is a collaborative platform focused on software development, its users are primarily technically proficient developers who not only use existing features but also frequently contribute to the improvement of software and tools. As a result, [A.3] Feature Suggestions is a common category in GitHub Issues. The [B] API category in the GitHub challenge taxonomy also has a significantly higher proportion compared to the OpenAI developer forum (i.e., 44.7% versus 22.9%). This further indicates that GitHub users are primarily highly skilled developers who are directly involved in software development, debugging, and maintenance, and thus more frequently engage with the underlying functionality and design of APIs. For these developers, the use, optimization, and error handling of APIs are central issues in their daily development tasks, leading to a substantial amount of API-related discussions on GitHub. Finally, issues on GitHub rarely discuss topics related to [E] GPT Builder, likely because Gemini and Llama do not support user plugin or GPT development in the same way as ChatGPT. This lack of support reduces the frequency of related issues. We also notice that some categories from the taxonomy constructed based on the OpenAI forum do not appear in this taxonomy. We speculate the reasons might be: certain issues are rarely discussed in project issues (e.g., [D.4] Promotion), advanced prompt design methods have not gained sufficient attention from developers (e.g., [F.2] Retrieval Augmented Generation to [F.4] In-context Learning), and the number of projects for specific applications may be limited (e.g., [C.6] Vision Capability).
7.2. Threats to Validity
@@ 7.2.1. Internal Validity
Expertise in constructing the challenge taxonomy: The threat is that insufficient expertise among annotators constructing the taxonomy could lead to an inaccurate reflection of the challenges developers face. To alleviate this threat, we require that those involved in the taxonomy development have relevant experience in LLM development and a foundational understanding of large language models. This ensures the taxonomy’s reliability and relevance.
Determining question difficulty levels: The threat is that the method used to determine the difficulty level of questions posed on the OpenAI forum may not accurately capture the actual difficulty of the questions. To alleviate this threat, we introduce additional metrics beyond the proportion of posts with an accepted solution, such as the average time to receive the first reply and the distribution of posts by the number of replies. This analysis from multiple perspectives provides a more comprehensive view to measure question difficulty.
@@
Manual labeling of topics: The threat is that manual labeling of topics could introduce subjective bias. To alleviate this threat, we employ a two-step labeling process: two independent annotators initially label the questions, and a third annotator, a domain expert, reviews these labels. Discrepancies are resolved through discussions among the annotators, thus reducing labeling bias to an acceptable level.
7.2.2. External Validity
Limitations of OpenAI LLMs: The threat is that focusing primarily on OpenAI’s LLMs in our empirical study may limit the generalizability of our findings to other LLM vendors. First, by analyzing the number of followers on GitHub accounts for LLMs, we found that ChatGPT has 86.6k followers, while Llama and Gemini have only 5.2k and 1k followers, respectively. Therefore, it can be inferred that the LLMs provided by OpenAI are currently more popular. Second, although different LLM vendors may offer varying features (such as image or audio processing capabilities), we believe that many of the challenges identified in our study (such as API usage, performance optimization, and security issues) are common across different LLMs. This implies that the insights from our study can inform development practices beyond OpenAI’s ecosystem. Finally, to demonstrate the generalizability of our constructed taxonomy, we further analyze challenges related to two other popular LLMs (such as Llama and Gemini) by mining issues from relevant projects on GitHub.
@@ 7.2.2. External Validity
Scope of data: An external challenge arises from the temporal limitation of our data, encompassing posts up to June 2024. As the LLM vendors continue to update their products, new challenges will likely emerge, which our current empirical study will not capture.
8. Related Work
8.1. LLM for Software Engineering
With the recent advancements in LLMs, there has been growing interest in applying these models to software engineering research, a field now referred to as LLM4SE (Hou et al., 2023; Fan et al., 2023). This area focuses on utilizing LLMs to enhance various software engineering activities. The primary activities that have garnered significant attention include software development, software quality assurance, and software maintenance. In this subsection, we summarize the recent research progress in LLM4SE based on these three core activities.
For the software development activity, the key tasks include code generation, code completion, and code summarization. Specifically, code generation involves automatically generating source code from task descriptions. For example, Yan et al. (Yan et al., 2023) explored ChatGPT’s ability to solve programming competition problems under zero-shot learning. Gu et al. (Gu et al., 2024) conducted an in-depth study of LLMs in domain-specific code generation, focusing on domains such as web, game, and math. To effectively incorporate domain knowledge, they utilized techniques like external knowledge inquirer, chain-of-thought prompting, and chain-of-thought fine-tuning. Liu et al. (Liu et al., 2024d, a) examined and detailed various issues related to the quality of code generated by ChatGPT. Liu et al. (Liu et al., 2024c) conducted a comprehensive evaluation of eight LLMs on five datasets to identify the challenges and limitations they face in different code generation scenarios. Code completion is the task of automatically suggesting the next portion of code based on the current context. For example, Eghbali et al. (Eghbali and Pradel, 2024) improved LLM-based code completion by retrieving suitable API references and iteratively querying the model with increasingly appropriate context information in the prompt. Tang et al. (Tang et al., 2023) incorporated domain-specific knowledge into LLMs by leveraging external information sources for code completion, without requiring fine-tuning of the underlying models. Li et al. (Li et al., 2023) used program structure-consistent mutations to test LLM-based code completion systems. These mutated inputs were then used to trigger potential errors in the code completion system. Code summarization is the task of automatically generating concise descriptions or explanations of code snippets. For example, Geng et al. (Geng et al., 2024) used LLMs to generate comments with multiple intents with in-context learning. Ahmed et al. (Ahmed et al., 2024) improved LLM-based code summarization by augmenting the prompt with semantic information. Sun et al. (Sun et al., 2024) conducted a comprehensive study on code summarization in the era of LLMs, examining various aspects such as prompt engineering techniques and temperature parameter settings to optimize performance. Zhao et al. (Zhao et al., 2024) improved LLM-based smart contract comment generation achieved through in-context learning.
For the software quality assurance activity, the key tasks include vulnerability detection, test generation, and bug localization. Specifically, vulnerability detection aims to identify security flaws or weaknesses in software code that could be exploited by malicious actors. For example, Lu et al. (Lu et al., 2023) assessed the effectiveness of prompt tuning for LLM-based vulnerability detection. Then they (Lu et al., 2024) improved LLM-based vulnerability detection with graph structure and in-contex learning. Steenhoek et al. (Steenhoek et al., 2024) leveraged a data flow analysis-inspired graph learning framework to improve LLM-based vulnerability detection. Chen et al. (Chen et al., 2023b) investigated the strengths and limitations of utilizing ChatGPT for detecting smart contract vulnerabilities. Test generation refers to the automatic generation of test cases or test scripts based on software requirements or existing code. For example, Schäfer et al. (Schäfer et al., 2023) examined the effectiveness of LLMs in generating unit tests, where the prompts of LLMs include the function’s signature and implementation, alongside usage examples extracted from the documentation. Yuan et al. (Yuan et al., 2023) performed a quantitative analysis and user study to systematically evaluate the quality of tests generated by ChatGPT, focusing on aspects such as correctness, sufficiency, readability, and usability. Ryan et al. (Ryan et al., 2024) proposed a code-aware prompting strategy for LLMs in test generation. Bug/fault localization refers to the process of automatically identifying the specific location or cause of a bug in the code. For example, Wu et al. (Wu et al., 2023) provided an in-depth investigation into ChatGPT’s capabilities for fault localization, considering the effects of prompt engineering and code context length. Kang et al. (Kang et al., 2024) proposed an LLM-based fault localization technique, which can generate both an explanation of the bug and a suggested fault location. Hossain et al. (Hossain et al., 2024) used LLMs to identify bug locations at the token level and then applied them to perform bug fixing.
For the software maintenance activity, the key tasks include program repair, code clone detection, and code review. Specifically, program repair refers to the automated process of identifying and fixing bugs/vulnerabilities in source code. For example, Jin et al. (Jin et al., 2023) enhanced LLM-based program repair by incorporating a static analyzer. Their approach involves a transformer encoder model, pretrained using contrastive learning, which is specifically designed to identify semantically similar bugs and corresponding fixes. Additionally, a decoder model is fine-tuned on supervised bug-fix data to generate effective repairs. Peng et al. (Peng et al., 2024) proposed a prompt-based approach for repairing Python type errors. This method begins by mining generalized fix templates, which serve as domain knowledge. These templates are then used to generate code prompts, guiding the model to produce accurate repairs for type errors. Code clone detection is the task of identifying identical or highly similar code fragments within or across software systems. For example, Dou et al. (Dou et al., 2023) conducted the first comprehensive evaluation of LLMs for code clone detection, examining various clone types, programming languages, and prompt strategies. Zhang et al. (Zhang and Saber, 2024) assessed the performance of ChatGPT for code clone detection. For example, they found that LLMs exhibited low effectiveness in detecting the most complex Type-4 code clones. Code review is a critical step in software quality assurance, involving tasks such as reviewer recommendation, code issue prediction, and review comment recommendation and generation. Li et al. (Li et al., 2022) proposed a pre-trained model designed specifically for code review tasks, such as code change quality estimation, review comment generation, and code refinement. Yu et al. (Yu et al., 2024) aimed to improve the accuracy and comprehensibility of code review by fine-tuning LLMs.
8.2. Studies on ML/DL Developers’ Challenges
@@
In traditional software development, developers face a broad range of challenges across different application domains such as mobile applications (Rosen and Shihab, 2016), web applications (Scoccia et al., 2021), and concurrency programs (Ahmed and Bagherzadeh, 2018). These challenges include understanding complex requirements, ensuring code quality and maintainability, handling bugs/security vulnerabilities, and integrating various software components.
The rapid development of machine learning (ML) technologies poses new challenges for software developers. Thung et al. (Thung et al., 2012) analyzed bug severity, bug fixing efforts, and bug impacts in ML systems. Alshangiti et al. (Alshangiti et al., 2019) demonstrated that questions in the process of ML application development are more difficult to answer than questions in other domains on Stack Overflow. These studies highlight the unique difficulties in developing and maintaining ML systems, such as data quality issues, model training complexities, and integrating ML models with traditional software components.
In the realm of deep learning (DL), DL developers face distinct challenges. For example, Zhang et al.(Zhang et al., 2019) found that program crashes, model migration, and implementation questions are the top three most frequently asked questions when developing DL applications. Morovati et al. (Morovati et al., 2024) investigated the challenges of deep reinforcement learning application development. Other researchers have characterized faults in software that make use of DL frameworks. For example, Zhang et al.(Zhang et al., 2018) categorized the symptoms and root causes of these DL bugs for TensorFlow and proposed strategies to detect and locate them. Then Islam et al.(Islam et al., 2019) and Humbatova et al.(Humbatova et al., 2020) extended the scope to include bugs in programs written based on more popular DL frameworks (such as Caffe, Keras, TensorFlow, Theano, and PyTorch), presenting more comprehensive results. Recently, Shen et al. also analyzed the DL compiler bugs (Shen et al., 2021; Chen et al., 2023a).
Empirical studies also highlighted specific issues, such as bugs in model optimization (Guan et al., 2023), compatibility problems (Wang et al., 2023; Guo et al., 2019), DL deployment challenges (Chen et al., 2020), and performance-related concerns (Cao et al., 2022).
In recent years, the rise of foundation models, particularly large language models, has led to remarkable progress across numerous domains, including software engineering. However, developers working with LLMs face a distinct set of challenges. Different from conventional ML/DL development, which often focuses on model training and optimization, LLM development introduces new issues, such as prompt design, the hallucination problem, result reproducibility, and API call costs. These challenges, combined with the multi-modal nature of LLMs (e.g., handling text, images, and audio), create complexities that are not typically encountered in traditional ML/DL workflows. Unfortunately, due to the rapid pace of development in this field, there are currently no comprehensive studies addressing these emerging challenges or offering clear guidelines for LLM developers. This study aims to fill this gap by conducting the first large-scale empirical study into the specific challenges faced by LLM application developers. By manually analyzing 2,364 questions from the OpenAI developer forum, this study constructs a taxonomy of LLM-related challenges and provides actionable insights for developing LLM-driven systems.
9. Conclusion
@@
In this study, we conduct a comprehensive analysis of the challenges faced by LLM developers. By examining relevant posts on the OpenAI developer forum, we observe that LLM-related development is gaining significant traction, with developers encountering numerous challenging issues compared to traditional software development. Our goal is to analyze the underlying challenges reflected in these questions. We investigate the popularity trends on the LLM developer forum and the difficulty levels of the problems raised by developers. Subsequently, we manually inspect 2,364 sampled posts related to LLM development and construct a detailed taxonomy consisting of 6 main categories and 26 subcategories, representing the challenges that LLM developers encounter. Finally, we discuss the implications of our findings for various stakeholders, including LLM developers and LLM vendors.
@@
For future research, there are several avenues to explore. First, expanding the analysis to include posts from other platforms, such as Stack Overflow, can provide a more comprehensive view of the challenges faced by LLM developers. Second, an in-depth study of the solutions proposed in response to the identified challenges can yield valuable insights into effective strategies and best practices. Finally, automating the categorization process using machine learning techniques can improve the efficiency and accuracy of the analysis.